# Quantum Hermitian conjugate and encoding unnormalized matrices

### v2

@@ II Extensions of matrix encoding technique
The matrix encoding implemented in the algorithms for matrix manipulations proposed in Ref.ZQKW_2024 (see also appendix in (ZBQKW_arXive2024 ) )uses two quantum subsystems $R$ and $C$ whose basis states label the rows and columns associated with each element $a_{jk}$ (complex in general) of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots,N-1\}$. Thus, the matrix $A$ is encoded as follows:
@@
We introduce two extensions to the algorithms of matrix manipulations developed in ZQKW_2024 . First of them is aimed on effective operations with complex matrices, Sec.II.1. The second one allows to reduce the constraint (2) imposed on the matrix elements, Sec.II.2. We deal with the square $N\times N$ matrices for convenience. If the original matrix is rectangular we turn it to the square matrix by adding the appropriate number of either zero rows or zero columns. We also require $N=2^{n}$, $n$ is a positive integer.
II.1 Hermitian conjugate
@@
where $SWAP_{RC}$ exchanges states of the subsystems $R$ and $C$ (in other words, it performs transposition of $A$), and $\sigma^{(z)}_{M}$ is the $\sigma^{(z)}$-operator which is applied to $M$ and thus performs complex conjugation of $A$. Thus the matrix encoded into the state $W|\Psi_{A}\rangle$ is the Hermitian conjugate of $A$,
$|\tilde{\Psi}_{A}\rangle=W|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{jkm}(-1)^{m}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (6)
The Hermitian conjugation algorithm can be used as a subroutine in other algorithms, as shown in Sec.III.1. The circuit for Hermitian conjugation is presented in Fig.1.
Figure 1: The circuit for Hermitian conjugation, $Z\equiv\sigma^{(z)}$.
We shall notice that using the new encoding of the complex matrix in the superposition state of the triple system $R\cup C\cup M$ requires appropriate modifications for the algorithms in Refs.ZQKW_2024 ; ZBQKW_arXive2024 . In addition, since algorithms of determinant calculation and matrix inversion ZBQKW_arXive2024 use the row-wise matrix encoding, each row requires its own additional one-qubit subsystem to label the real and imaginary parts of its matrix elements, i.e., $N$ one-qubit subsystems $M_{i}$, $i=0,\dots,N-1$, are required. In Sec.III, we consider modification of multiplication algorithm.
II.2 Weakened normalization constraint for matrix encoding
Equality (2) represents a constraint imposed on the absolute values of matrix elements. However, we can weaken this constraint reducing it to the inequality
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}<1.$ (7)
To satisfy inequality (7), the matrix $A$ has to be properly prepared. Namely, instead of the matrix $A$ we encode the matrix $\tilde{A}=\{\tilde{a}_{ij}:i,j=0,\dots,N-1\}$,
$\tilde{A}=\frac{1}{\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}+c}A,\;\;c>0,$ (8)
where the positive parameter $c$ is conventional, it is required to guarantee that condition (7) for the matrix $\tilde{A}$ is a strict inequality. To encode the prepared matrix $\tilde{A}$, we introduce, along with the subsystems $R$ and $C$ enumerating rows and columns of $\tilde{A}$, the one-qubit subsystem $K$ and replace the initial state (1) with the following one:
$|\Psi_{A}\rangle$ $=$ $b|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\tilde{a}_{jk}|j\rangle_{R}|k\rangle_{C}|1\rangle_{K}$ (9)
@@ II.2 Weakened normalization constraint for matrix encoding
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}=1,$ (10)
where $b$ serves to satisfy the constraint (10) and can be determined as
$|b|^{2}=1-\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}.$ (11)
II.3 Circuit for new encoding of initial state
Now we combine the modifications of the state encoding presented in Eqs.(4) and (9) and encode the elements $\tilde{a}_{ij}$ decomposed into the real and imaginary parts (3) as follows:
$|\Psi_{A}\rangle$ $=$ $(b^{(i)}_{0}|0\rangle_{M}+b^{(i)}_{1}|1\rangle_{M})|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+$ (13)
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}(\tilde{a}_{jk0}|0\rangle_{M}+\tilde{a}_{jk1}|1\rangle_{M})|j\rangle_{R}|k\rangle_{C}|1\rangle_{K},$
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|\tilde{a}_{jk}|^{2}=1,$
Replacing matrix representation (1) with representation (13) is shown in Fig.2.
Figure 2: Replacement of matrix encoding (1) with encoding (13) for organizing effective manipulations with complex matrices and weakening the constraints on the matrix elements.
Below we omit tilde over the elements of the encoded matrix for the brevity, i.e., write $a_{jk}$ instead of $\tilde{a}_{jk}$ assuming that elements $a_{jk}$ satisfy inequality (7).
III Matrix multiplication
We modify algorithm for matrix multiplication proposed in ZQKW_2024 and improved in ZBQKW_arXive2024 . Modification is based on passing from encoding (1) (used in ZQKW_2024 ) to encoding (13) for the square $N\times N$ matrices $A^{(1)}$ and $A^{(2)}$ ($A^{(i)}=\{a^{(i)}_{jk}:j,k=0,\dots,N-1\}$, $i=1,2$) according to Sec.II. We assume that the both matrices $A^{(1)}$ and $A^{(2)}$ are prepared in accordance with formula (8) and therefore their elements satisfy inequality (7).
We first introduce four $n$-qubit registers $R^{(i)}$ and $C^{(i)}$ , $i=1,2$, enumerating, respectively, the rows and columns of the matrices $A^{(i)}$, $1,2$. Then, we introduce four one-qubit subsystems $M^{(i)}$ and $K^{(i)}$, $i=1,2$. The subsystems $M^{(i)}$, $i=1,2$, are used to label the real and imaginary parts of the matrix elements of the $i$th matrix ($a^{(i)}_{jk}=a^{(i)}_{jk0}+ia^{(i)}_{jk1}$), and also to label the real and imaginary parts of the parameters $b^{(i)}$ ($b^{(i)}=b^{(i)}_{0}+ib^{(i)}_{1}$). The subsystems $K^{(i)}$, $i=1,2$, are needed to weaken the constraints imposed on the elements $a^{(i)}_{jk}$. Thus, the pure states encoding the elements of matrices $A^{(i)}$, $i=1,2$, are in form (13) with normalization (13):
$|\Psi^{(i)}\rangle$ $=$ $(b^{(i)}_{0}|0\rangle_{M^{(i)}}+b^{(i)}_{1}|1\rangle_{M^{(i)}})|0\rangle_{R^{(i)}}|0\rangle_{C^{(i)}}|0\rangle_{K^{(i)}}+$ (15)
$\sum_{j,k=0}^{N-1}(a^{(i)}_{jk0}|0\rangle_{M^{(i)}}+a^{(i)}_{jk1}|1\rangle_{M^{(i)}})|j\rangle_{R^{(i)}}|k\rangle_{C^{(i)}}|1\rangle_{K^{(i)}},$
$|b^{(i)}|^{2}+\sum_{j,k}|a^{(i)}_{jk}|^{2}=1,$
where the constant parameters $b^{(i)}$ are introduced in accordance with Sec.II.2. The initial state of the whole system reads
$|\Phi_{0}\rangle=|\Psi_{2}\rangle\otimes|\Psi_{1}\rangle=$ (16)
@@ III Matrix multiplication
×$\times|j_{1}\rangle_{R_{1}}|k_{1}\rangle_{C^{(1)}}|j_{2}\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{0}\rangle,$
@@
From the superposition state $|\Phi_{0}\rangle$, we have to select the terms with the states $|k\rangle_{C^{(1)}}|k\rangle_{R^{(2)}}$, $k=0,\dots,N-1$, i.e., $k_{1}=j_{2}$ in the second part in the rhs of formula (16). For that purpose we apply the CNOTs $W_{j}$ to the $j$th qubits of $C^{(1)}$ and $R^{(2)}$,
@@
where $\sigma^{(x)}_{R^{(2)}_{j}}$ and $I_{R^{(2)}_{j}}$ are, respectively, the $\sigma^{(x)}$-operator and identity operator applied to the $j$th qubit of the subsystem $R^{(2)}$. We have chosen the qubits of $C^{(1)}$ as control ones. All operators $W_{j}$, $j=1,\dots,n$, commute with each other. Applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$,
@@
×$\times|j_{1}\rangle_{R^{(1)}}|j\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{1}\rangle.$
The second part in the rhs of (19) collets the terms with the state $|0\rangle_{R_{2}}$ (the other terms are garbage by definition; their are collected in $|g_{1}\rangle$ above). We note that the first part in the rhs of (19) remains unchanged after applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$.
@@
to the qubits of $C^{(1)}$ and select the terms with $|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}$ putting others to the garbage $|g_{2}\rangle$:
@@
×$\times|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{2}\rangle.$
We note that the operator $W^{(1)}_{C^{(1)}}$ transforms $|g_{1}\rangle$ as well. However, we don’t need details of that transformation. The similar statement can be referred to the operators $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ and $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$ below.
Now, to complete multiplication, we have to organize proper combinations of products of real and imaginary parts of matrices $A^{(i)}$, $i=1,2$, to form the real and imaginary parts of the result. To this end we introduce the following control operator:
@@
where $H_{M^{(2)}}$ is the set of the Hadamard operators applied to single qubit of $M^{(2)}$, and apply $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ to $|\Phi_{2}\rangle$, thus obtaining
@@
×$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
@@
×$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+|g_{3}\rangle.$
Here the second term contains the desired matrix product and all extra terms are collected in the garbage $|g_{3}\rangle$. We emphasize that the operator $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ is a novelty in comparison with multiplication algorithm in ZBQKW_arXive2024 .
Next, to label and remove the garbage, we prepare two one-qubit ancillae $B$ and $\tilde{B}$ in the ground state $|0\rangle_{B}|0\rangle_{\tilde{B}}$, introduce the projector
@@
and the controlled operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}$,
$W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}=P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}\otimes\sigma^{(x)}_{B}\sigma^{(x)}_{\tilde{B}}$ (27)
$+(I_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}-P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}})\otimes I_{B\tilde{B}},$
of the depth $O(n)$ with $2(n+1)$-qubit control register. Applying this operator to $|\Phi_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}$ we obtain
$|\Phi_{4}\rangle=W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B\tilde{B}}|\Phi_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}=$ (28)
@@
×$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}|1\rangle_{B}|1\rangle_{\tilde{B}}+$
$\frac{1}{2^{(n+1)/2}}\sum_{{j_{1},k_{2}}\atop{j=0}}^{N-1}\Big{(}(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}\ |1\rangle_{B}|1\rangle_{\tilde{B}}+|g_{3}\rangle|0\rangle_{B}|0\rangle_{\tilde{B}}.$
Now we perform the conditional measurement FZQW_arxive2025
$W^{(4)}_{B\tilde{B}}=|1\rangle_{B}\,{}_{\tilde{B}}\langle 1|\otimes M_{\tilde{B}}+|0\rangle_{B}\,{}_{\tilde{B}}\langle 0|\otimes I_{\tilde{B}},$ (29)
where $M_{\tilde{B}}$ is the measurement operator applied to $\tilde{B}$. Then,
$|\Phi_{5}\rangle=W^{(4)}_{B\tilde{B}}|\Phi_{4}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B},$ (30)
$|\Psi_{out}\rangle=G^{{-1}}\Big{(}\Big{(}(b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big{)}$
×$\times|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K_{1}}+$
@@
×$\times|j_{1}\rangle_{R^{(1)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\Big{)}=$

### v4

@@ IV Extensions of matrix encoding technique
The matrix encoding implemented in the algorithm for matrix multiplication uses two $n$-qubit quantum subsystems $R$ and $C$ whose basis states label, respectively, the rows and columns associated with each element $a_{jk}$ (complex in general) of the $N\times N$ matrix $A=\{a_{jk}:j,k=0,\dots,N-1\}$, $N=2^{n}$. Thus, the matrix $A$ is encoded as follows:
@@
We introduce two extensions to the algorithms of matrix manipulations.The first of them is aimed on effective implementing the operations with complex matrices, Sec.IV.1. The second one allows to weaken the constraint, Eq.(43), imposed on the matrix elements by the normalization condition holding for pure quantum states, Sec.IV.2. We deal with the square $N\times N$ matrices for convenience. If the original matrix is rectangular we turn it to the square matrix by appending the appropriate number of either zero rows or zero columns. In Sec. IV.3, both above extensions are combined in the single circuit. Then, in Sec.V, both these extensions will be implemented in the matrix-multiplication algorithm. Therein, we also compare the results of application of the usual and controlled measurements.
IV.1 Hermitian conjugate
@@
where $SWAP_{RC}$ exchanges states of the subsystems $R$ and $C$ (in other words, it performs transposition of $A$), and $\sigma^{(z)}_{M}$ is the $\sigma^{(z)}$-operator applied to $M$. This operator changes the sign of the excited state $|1\rangle_{M}$ and thus performs the complex conjugation of $A$. Therefore, the matrix encoded into the state $W|\Psi_{A}\rangle$ is the Hermitian conjugate of $A$,
$|\tilde{\Psi}_{A}\rangle=W|\Psi_{A}\rangle=\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}\sum_{m=0}^{1}a_{kjm}(-1)^{m}|j\rangle_{R}|k\rangle_{C}|m\rangle_{M}.$ (47)
The Hermitian conjugation algorithm can be used as a subroutine in other algorithms, as shown in Sec.V.1. The circuit for the Hermitian conjugation is presented in Fig.3.
Figure 3: The circuit for the Hermitian conjugation, $Z\equiv\sigma^{(z)}$.
We shall notice that using the new encoding of the complex matrix in the superposition state of the triple system $R\otimes C\otimes M$ requires appropriate modifications for the multiplication algorithm which is considered in Sec.V. In addition, since the algorithms for determinant calculation and matrix inversion ZBQKW_arXive2024 use the row-wise matrix encoding, each row requires its own additional one-qubit subsystem to label the real and imaginary parts of its matrix elements, i.e., $N$ one-qubit subsystems $M_{i}$, $i=0,\dots,N-1$, are required for those algorithms.
IV.2 Weakened normalization constraint for matrix encoding
Eq.(43) represents a constraint imposed on the modulus of the matrix elements. However, we can weaken this constraint reducing it to the inequality
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}\leq 1.$ (48)
This can be done if we introduce, along with the subsystems $R$ and $C$ enumerating rows and columns of $A$, the one-qubit subsystem $K$ and replace the initial state given in Eq.(42) with the following one:
$|\Psi_{A}\rangle$ $=$ $b|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}a_{jk}|j\rangle_{R}|k\rangle_{C}|1\rangle_{K}$ (49)
@@ IV.2 Weakened normalization constraint for matrix encoding
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$ (50)
Eq.(50) yields inequality (48). Thus, replacing Eq.(43) with inequality (48) becomes possible due to introducing the additional probability amplitude $b$ into the initial state, Eq.(49). Again, to weaken the normalization constraint in the row-wise encoding ZBQKW_arXive2024 , an additional qubit is requored for encoding each row.
IV.3 Circuit for new encoding of initial state
Now we combine the modifications of the state encoding presented in Eqs.(45) and (49) and encode the elements $a_{ij}$ decomposed into the real and imaginary parts, shown in Eq.(44), as follows:
$|\Psi_{A}\rangle$ $=$ $(b_{0}|0\rangle_{M}+b_{1}|1\rangle_{M})|0\rangle_{R}|0\rangle_{C}|0\rangle_{K}+$ (52)
$\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}(a_{jk0}|0\rangle_{M}+a_{jk1}|1\rangle_{M})|j\rangle_{R}|k\rangle_{C}|1\rangle_{K},\;\;b=b_{0}+ib_{1},$
$|b|^{2}+\sum_{j=0}^{N-1}\sum_{k=0}^{N-1}|a_{jk}|^{2}=1.$
Replacement of matrix encoding via Eq. (42) with encoding via Eq.(52) is shown in Fig.4.
Figure 4: Replacement of matrix encoding via Eq.(42) with encoding via Eq.(52) for organizing effective manipulations with complex matrices and weakening the constraint on the matrix elements.
V Matrix multiplication
In this section we propose the matrix multiplication algorithm ZQKW_2024 ; ZBQKW_arXive2024 that takes into account the extensions discussed in Secs. IV, as well as implements the controlled measurement described in Sec.II.2.
Proposed modification is based on passing from encoding via Eq.(42) (used in ZQKW_2024 ) to encoding via Eq.(52) for the square $N\times N$ matrices $A^{(1)}$ and $A^{(2)}$ ($A^{(l)}=\{a^{(l)}_{jk}:j,k=0,\dots,N-1\}$, $l=1,2$) according to Sec.IV.
[…]
$|\Psi^{(l)}\rangle$ $=$ $(b^{(l)}_{0}|0\rangle_{M^{(l)}}+b^{(l)}_{1}|1\rangle_{M^{(l)}})|0\rangle_{R^{(l)}}|0\rangle_{C^{(l)}}|0\rangle_{K^{(l)}}+$ (54)
$\sum_{j,k=0}^{N-1}(a^{(l)}_{jk0}|0\rangle_{M^{(l)}}+a^{(l)}_{jk1}|1\rangle_{M^{(l)}})|j\rangle_{R^{(l)}}|k\rangle_{C^{(l)}}|1\rangle_{K^{(l)}},$
$|b^{(l)}|^{2}+\sum_{j,k}|a^{(l)}_{jk}|^{2}=1,\;\;l=1,2,$
where the constant parameters $b^{(l)}$ are introduced in accordance with Sec.IV.2. The initial state of the whole system reads
$|\Phi_{0}\rangle=|\Psi^{(2)}\rangle\otimes|\Psi^{(1)}\rangle=$ (55)
@@ V Matrix multiplication
$\times|j_{1}\rangle_{R_{1}}|k_{1}\rangle_{C^{(1)}}|j_{2}\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{0}\rangle,$
@@
From the superposition state $|\Phi_{0}\rangle$, we have to select the terms with the states $|k\rangle_{C^{(1)}}|k\rangle_{R^{(2)}}$, $k=0,\dots,N-1$, i.e., $k_{1}=j_{2}$ in the second part in the right-hand side of Eq.(55). For that purpose we apply the CNOTs $W_{j}$ to the $j$th qubits of $C^{(1)}$ and $R^{(2)}$,
@@
where $\sigma^{(x)}_{R^{(2)}_{j}}$ and $I_{R^{(2)}_{j}}$ are, respectively, the $\sigma^{(x)}$-operator and identity operator applied to the $j$th qubit of the subsystem $R^{(2)}$. We have chosen the qubits of $C^{(1)}$ as controlling ones. All operators $W_{j}$, $j=1,\dots,n$, commute with each other. Applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$,
@@
$\times|j_{1}\rangle_{R^{(1)}}|j\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{1}\rangle.$
The first and second parts in the right-hand side of Eq.(V) collect the terms with the state $|0\rangle_{R_{2}}|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, (the other terms are garbage by definition, they are collected in $|g_{1}\rangle$ above). We note that the first part in the right-hand side of Eq.(V) remains unchanged after applying the operator $W^{(0)}_{C^{(1)}R^{(2)}}$. We also note that the operator $W^{(0)}_{C^{(1)}R^{(2)}}$ transforms $|g_{0}\rangle$ as well. However, we don’t need details of that transformation. The similar statement can be referred to the operators $W^{(1)}_{C^{(1)}}$, $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ and $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}$ below.
@@
to the qubits of $C^{(1)}$ and select the terms with $|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|j\rangle_{K^{(1)}}|j\rangle_{K^{(2)}}$, $j=0,1$, putting others to the garbage $|g_{2}\rangle$:
@@
$\times|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|1\rangle_{K^{(2)}}+|g_{2}\rangle.$
Now, to complete multiplication, we have to organize proper combinations of products of real and imaginary parts of matrices $A^{(i)}$, $i=1,2$, to form the real and imaginary parts of the result. To this end we introduce the following controlled operator:
@@
where $H_{M^{(2)}}$ is the Hadamard operator applied to the single qubit of $M^{(2)}$, and apply $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ to $|\Phi_{2}\rangle$, thus obtaining
@@
$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}$
@@
$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}+|g_{3}\rangle.$
Here the first term plays an auxiliary role similar to the first term of Eq.(54), while the second term contains the desired matrix product. All extra terms are collected in the garbage $|g_{3}\rangle$. We emphasize that the operator $W^{(2)}_{M^{(1)}M^{(2)}K^{(1)}K^{(2)}}$ is a novelty in comparison with multiplication algorithm in ZBQKW_arXive2024 .
Next, to label and remove the garbage, we prepare the two-qubit ancillae $B$ in the ground state $|0\rangle_{B}$, introduce the projector
@@
and the controlled operator $W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}$,
$W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}=P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}\otimes\sigma^{(x)}_{B_{1}}\sigma^{(x)}_{B_{2}}$ (66)
$+(I_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}}-P_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}})\otimes I_{B},$
of the depth $O(n)$ with $2(n+1)$-qubit control register. Here $B_{i}$, $i=1,2$, means the $i$th qubit of the ancilla $B$. Applying this operator to $|\Phi_{3}\rangle|0\rangle_{B}$ we obtain
$|\Phi_{4}\rangle=W^{(3)}_{C^{(1)}R^{(2)}M^{(2)}K^{(2)}B}|\Phi_{3}\rangle|0\rangle_{B}=$ (67)
@@
$\times|0\rangle_{M^{(2)}}|0\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|0\rangle_{C^{(2)}}|0\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}|1\rangle_{B_{1}}|1\rangle_{B_{2}}+$
$\frac{1}{2^{(n+1)/2}}\sum_{{j_{1},k_{2},}\atop{j=0}}^{N-1}\Big((a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}0}-a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}1})|0\rangle_{M^{(1)}}+(a^{(1)}_{j_{1}j0}a^{(2)}_{jk_{2}1}+a^{(1)}_{j_{1}j1}a^{(2)}_{jk_{2}0})|1\rangle_{M^{(1)}}\Big)$
$\times|0\rangle_{M^{(2)}}|j_{1}\rangle_{R^{(1)}}|0\rangle_{C^{(1)}}|0\rangle_{R^{(2)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}|0\rangle_{K^{(2)}}\ |1\rangle_{B_{1}}|1\rangle_{B_{2}}+|g_{3}\rangle|0\rangle_{B_{1}}|0\rangle_{B_{2}}.$
Now we have to remove the garbage marked by the state $|0\rangle_{B_{1}}|0\rangle_{B_{2}}$. Before implementing the controlled measurement described in Sec.II.2.2, we apply the most straightforward way of garbage removal that is the usual measurement of the state of $B_{2}$ with the desired output $|1\rangle_{B_{2}}$. Of course, this way requires multiple running of the algorithm because the single run with the subsequent measurement might yield undesired result $|0\rangle_{B_{2}}$ that must be disregarded. But once the desired output is obtained, the quantum state of the remaining system becomes
$|\Phi_{5}\rangle=|\Psi_{out}\rangle\,|0\rangle_{M_{2}}|0\rangle_{C_{1}}|0\rangle_{R_{2}}|0\rangle_{K_{2}}|1\rangle_{B_{1}},$ (68)
$|\Psi_{out}\rangle=G^{{-1}}\Big(\Big((b^{(1)}_{0}b^{(2)}_{0}-b^{(1)}_{1}b^{(2)}_{1})|0\rangle_{M^{(1)}}+(b^{(1)}_{0}b^{(2)}_{1}+b^{(1)}_{1}b^{(2)}_{0})|1\rangle_{M^{(1)}}\Big)$ (69)
$\times|0\rangle_{R^{(1)}}|0\rangle_{C^{(2)}}|0\rangle_{K_{1}}+$
@@
$\times|j_{1}\rangle_{R^{(1)}}|k_{2}\rangle_{C^{(2)}}|1\rangle_{K^{(1)}}\Big)=$